Lemuel bought bottles of food supplements at Php 1 800 each. He is planning to sell this at 15% markup on cost. At this selling price and a fixed cost of Php 40 000, how many bottles of food supplement must he sell to be able to breakeven? If he wants to earn a minimum of Php 100 000 by just selling these bottles, how many bottles must he sell?


1
Expert's answer
2022-05-19T19:11:50-0400

Solution:


Assume that he sells x bottles of food supplement:



"1800x+40000 = 1800(1+15\% ) x""270 x = 40000""x=148"

Earn Php 100000:



"1800x+4000+100000=1800(1+15\%)x""270x=140000""x=518"